OpenStudy (nwolf):

Anyone feel like helping me out with this problem?? The percent (in decimal form) of battery power remaining x hours after you turn on a laptop computer is y = -0.25x + 1. Show your work and explain your thinking. a. What is the x-intercept of this equation? b. Explain the meaning of this number in real life terms. c. What is the y-intercept of this equation? d. Explain the meaning of this number in real life terms

OpenStudy (anonymous):

x = 4 is the x intercept. It means the number of hours it will take for the power to be dropped to zero. y = 1 is the y intercept. It means that the batter power left when you first turn on laptop.

OpenStudy (nwolf):

do you mind explaining to me how you were able to figure out the problem

OpenStudy (anonymous):

Well.....let us start with the equation: y = 0.25x + 1 where y is "power" and x is "hours". X intercept is where y = 0. So, set y=0. You get 0.25x+1 = 0 => x = -1/0.25 = -4 (sorry, I said 4 but it is a negative because time elapses). That means, 4 hours have to elapse for power to go from 100% level to ZERO level. Y intercept is where x = 0. So, set x = 0. You get y = 1. This means that at the beginning the power is 100%.

OpenStudy (anonymous):